# Reward Modeling from Natural Language Human Feedback

### v1

@@ 1 Introduction
Recently, generative reward models (GRMs) have emerged as a powerful solution to enhance the general capabilities of large language models (LLMs) (Liu et al., 2025; Wang et al., 2025b; Yuan et al., 2024; Guo et al., 2025b; Huang et al., 2025; Xu et al., 2025a). Unlike traditional scalar reward models that output a single numerical score, GRMs generate natural language rationales with detailed critiques before making judgments, leading to significantly higher robustness, generalizability, and accuracy through explicit reasoning (Mahan et al., 2024; Malik et al., 2025; Tan et al., 2024; Liu et al., 2024b; Wang et al., 2025b; Ye et al., 2024b; Zhang et al., 2025b; Shao et al., 2025b). Compared to rule-based verifiers that are restricted to verifiable tasks such as mathematical problems (Xie et al., 2025; Shao et al., 2024; Li et al., 2025c; Yu et al., 2025b; Li et al., 2025d) or code execution (Seed et al., 2025; Hui et al., 2024; Wang et al., 2025a), GRMs can effectively evaluate a much broader range of tasks including social intelligence (Yu et al., 2025a; Zhang et al., 2025a; Li et al., 2025a; Mao and Tao, 2025; Zhou et al., 2025a; Zhang et al., 2025a), roleplay (Zhou et al., 2025b; Lu et al., 2024; 2025; Qin et al., 2025; Nath et al., 2025), and image generation (Zhang et al., 2025c; Wu et al., 2025; Zhou et al., 2025c; Yang et al., 2025a), etc. Typically in a pairwise reward setting, GRMs produce a detailed CoT which mainly serves to derive effective critiques [^1] In this paper, we use “process” to refers to these critiques, instead of full reasoning process. of two responses, followed by an explicit outcome indicating which response is superior (Yuan et al., 2024; Guo et al., 2025b; Huang et al., 2025; Xu et al., 2025a).
@@
The current GRM training in reinforcement learning with verifiable reward (RLVR) typically relies on outcome-only supervision, primarily through binary pairwise preference labels (Wang et al., 2025b; Yuan et al., 2024; Guo et al., 2025b). However, we argue that while outcome reward are effective for mathematical tasks, they are fundamentally inadequate for GRM tasks. This disparity arises from the nature of their respective solution spaces: mathematical tasks typically have large solution spaces where a correct outcome strongly implies correct reasoning trajectory. In contrast, GRM tasks operate within a constrained binary solution space with only two possible outcomes (preference for response A or B). Such a binary classification task makes GRMs susceptible to guessing the correct preference label without generating sound critiques, e.g., objectively incorrect or nitpicky (non-essential) critiques (examples in Figure 1). We illustrate the mismatch in the proportions of process and outcome supervision between mathematical tasks and GRM tasks in Figure 2. Sometimes, mathematical tasks also exhibit small solution spaces (e.g., multiple-choice questions or true/false problems), and practitioners typically address this by either removing such data (Team et al., 2025; Cui et al., 2025) or reformulating them as fill-in-the-blank questions to expand the solution space (Albalak et al., 2025). However, in GRM tasks, the pairwise comparison formulation is inherently essential and cannot be abandoned or restructured. As a result, given the extremely restricted solution space of binary tasks, these spurious successes introduce substantial noise into the reward signal when using RL to train GRMs, thereby impairing the learning effectiveness. Specifically, the outcome-process misalignment resulting from outcome-only supervision might lead the model to exploit spurious correlations rather than developing genuine reasoning capabilities, causing the policy to converge toward producing erroneous critiques and thereby limiting the generalization and advancement potential of GRMs’ training.
@@
Motivated by these observations, we propose incorporating the similarity between human critiques and GRM-generated critiques as an additional process reward term alongside the outcome reward during GRM training. This approach yields significant improvements in both critique quality and outcome accuracy (Section 3.1). However, human critique annotation faces critical scalability challenges due to high costs and the difficulty of obtaining high-quality annotations (Wang et al., 2025b). In practice, existing data resources rarely include human critiques, with most datasets containing only outcome labels (Lai et al., 2024; Han et al., 2024; Wang et al., 2024c; Liu et al., 2024a). To address this limitation, we propose using MetaRM, an auxiliary meta reward model that learns to predict process reward from training data with human critiques and generalizes to unlabeled data with only outcome labels (Section 3.2.2). Additionally, to handle distribution shift as the policy evolves during RL training, we introduce an online MetaRM framework that continuously updates the MetaRM alongside GRM updates, enabling more effective process supervision at scale (Section 3.2.3). Our experiments demonstrate that online MetaRM-based training achieves comparable performance to full human critique supervision while significantly reducing annotation requirements (Section 3.2.4).
@@ 2.2 Outcome–Process Inconsistency
To validate the outcome-process inconsistency, we conduct experiments on two distinct tasks: mathematical reasoning on MATH-500 (Hendrycks et al., 2021) and pairwise rewarding on validation set of HelpSteer3 (Wang et al., 2025b). For mathematical reasoning, we use gemini-2.5-pro to compare model-generated solutions against ground-truth solutions, assessing both process validity and answer correctness. For pairwise rewarding, we evaluate outcome accuracy using a rule-based verifier and process accuracy by computing the similarity between model-generated critiques and human critiques using gemini-2.5-pro (whose accuracy we validate in Section 2.3).
@@ 4.3 Ablation Study
4.4 Performance of Downstream Tasks
Table 4: Evaluation results on downstream tasks.
@@ G.1 Benchmarks
RM-Bench (Liu et al., 2024b) 2024.10 Chat, Safety, Code, Math Verifiable correctness, deliberate errors
@@
SCAN-HPD (Wang et al., 2025c) 2025.05 Writing, Roleplay, Knowledge, Coding, Math, Reasoning Majority human voting
HelpSteer3 (Wang et al., 2025b) 2025.05 General chat, STEM, Code, Multilingual Majority human voting
@@ G.2 Training Datasets
HelpSteer3 (Wang et al., 2025b) 36,255 General Domain
@@ H.1 Ablation Study on MetaRM
Appendix I Prompt

### v2

@@ 1 Introduction
Recently, generative reward models (GRMs) have emerged as a powerful solution to enhance the general capabilities of large language models (LLMs) (Liu et al., 2025; Wang et al., 2025c; Yuan et al., 2024; Guo et al., 2025b; Huang et al., 2025; Xu et al., 2025a). Unlike traditional scalar reward models that output a single numerical score, GRMs generate natural language rationales with detailed critiques before making judgments, leading to significantly higher robustness, generalizability, and accuracy through explicit reasoning (Mahan et al., 2024; Malik et al., 2025; Tan et al., 2024; Liu et al., 2024c; Wang et al., 2025c; Ye et al., 2024b; Zhang et al., 2025b; Shao et al., 2025b). Compared to rule-based verifiers that are restricted to verifiable tasks such as mathematical problems (Xie et al., 2025; Shao et al., 2024; Li et al., 2025c; Yu et al., 2025b; Li et al., 2025d) or code execution (Seed et al., 2025; Hui et al., 2024; Wang et al., 2025a), GRMs can effectively evaluate a much broader range of tasks including social intelligence (Yu et al., 2025a; Zhang et al., 2025a; Li et al., 2025a; Mao and Tao, 2025; Zhou et al., 2025a; Zhang et al., 2025a), roleplay (Zhou et al., 2025b; Lu et al., 2024; 2025; Qin et al., 2025; Nath et al., 2025), and image generation (Zhang et al., 2025c; Wu et al., 2025; Zhou et al., 2025c; Yang et al., 2025a), etc. Typically in a pairwise reward setting, GRMs produce a detailed CoT which mainly serves to derive effective critiques [^1] In this paper, we use “process” to refers to these critiques, instead of full reasoning process. of two responses, followed by an explicit outcome indicating which response is superior (Yuan et al., 2024; Guo et al., 2025b; Huang et al., 2025; Xu et al., 2025a).
@@
The current GRM training in reinforcement learning with verifiable reward (RLVR) typically relies on outcome-only supervision, primarily through binary pairwise preference labels (Wang et al., 2025c; Yuan et al., 2024; Guo et al., 2025b). However, we argue that while outcome reward are effective for mathematical tasks, they are fundamentally inadequate for GRM tasks. This disparity arises from the nature of their respective solution spaces: mathematical tasks typically have large solution spaces where a correct outcome strongly implies correct reasoning trajectory. In contrast, GRM tasks operate within a constrained binary solution space with only two possible outcomes (preference for response A or B). Such a binary classification task makes GRMs susceptible to guessing the correct preference label without generating sound critiques, e.g., objectively incorrect or nitpicky (non-essential) critiques (examples in Figure 1). We illustrate the mismatch in the proportions of process and outcome supervision between mathematical tasks and GRM tasks in Figure 2. Sometimes, mathematical tasks also exhibit small solution spaces (e.g., multiple-choice questions or true/false problems), and practitioners typically address this by either removing such data (Team et al., 2025; Cui et al., 2025) or reformulating them as fill-in-the-blank questions to expand the solution space (Albalak et al., 2025). However, in GRM tasks, the pairwise comparison formulation is inherently essential and cannot be abandoned or restructured. As a result, given the extremely restricted solution space of binary tasks, these spurious successes introduce substantial noise into the reward signal when using RL to train GRMs, thereby impairing the learning effectiveness. Specifically, the outcome-process misalignment resulting from outcome-only supervision might lead the model to exploit spurious correlations rather than developing genuine reasoning capabilities, causing the policy to converge toward producing erroneous critiques and thereby limiting the generalization and advancement potential of GRMs’ training.
@@
Motivated by these observations, we propose incorporating the similarity between human critiques and GRM-generated critiques as an additional process reward term alongside the outcome reward during GRM training. This approach yields significant improvements in both critique quality and outcome accuracy (Section 3.1). However, human critique annotation faces critical scalability challenges due to high costs and the difficulty of obtaining high-quality annotations (Wang et al., 2025c). In practice, existing data resources rarely include human critiques, with most datasets containing only outcome labels (Lai et al., 2024; Han et al., 2024; Wang et al., 2024c; Liu et al., 2024a). To address this limitation, we propose using MetaRM, an auxiliary meta reward model that learns to predict process reward from training data with human critiques and generalizes to unlabeled data with only outcome labels (Section 3.2.2). Additionally, to handle distribution shift as the policy evolves during RL training, we introduce an online MetaRM framework that continuously updates the MetaRM alongside GRM updates, enabling more effective process supervision at scale (Section 3.2.3). Our experiments demonstrate that online MetaRM-based training achieves comparable performance to full human critique supervision while significantly reducing annotation requirements (Section 3.2.4).
@@
  • We explain four reward modeling mechanisms from theoretical perspective, clarifying the evolution of reward modeling and the necessity of our method, then identify the theory-practice gaps remain (Appendix I).
@@ 2.2 Outcome–Process Inconsistency
To validate the outcome-process inconsistency, we conduct experiments on two distinct tasks: mathematical reasoning on MATH-500 (Hendrycks et al., 2021) and pairwise rewarding on validation set of HelpSteer3 (Wang et al., 2025c). For mathematical reasoning, we use gemini-2.5-pro to compare model-generated solutions against ground-truth solutions, assessing both process validity and answer correctness. For pairwise rewarding, we evaluate outcome accuracy using a rule-based verifier and process accuracy by computing the similarity between model-generated critiques and human critiques using gemini-2.5-pro (whose accuracy we validate in Section 2.3).
@@ 4.3 Ablation Study
4.4 Performance of Downstream Tasks through Reinforcement Learning
In this section, we adopt RM-NLHF-Qwen-7B and Outcome-Only-Qwen-7B as the reward models to perform RL training on DeepSeek-Distilled-Qwen-7B. We use prompts from HelpSteer3 as the training set and apply GRPO for 100 training steps. Since both reward models are pairwise in nature, we convert them into pointwise scores via Double-Elimination Tournament (Zhang et al., 2026).
Results presented in Table 4 demonstrate that our method holds a substantial advantage over the baseline. Across a diverse set of SOTA judge models (gemini-3-pro, gpt-5.1, qwen3-max-thinking and kimi-k2.5 (Team et al., 2026)), our approach outperforms the baseline by an average margin of 12.92%.
Since GRM is known to exhibit a length bias (Dubois et al., 2024), we further analyze the output length throughout training, as shown in Figure 7. The model trained with RM-NLHF actually produces shorter responses than the baseline, confirming that the gains stem from improved quality rather than verbosity.
Table 4: Win rates (%) of RL-trained models using RM-NLHF vs. Outcome-Only as the reward model, evaluated on Arena-Hard-V2.0 (Li et al., 2025b) with different judge models.
Judge Model
gemini-3-pro gpt-5.1 qwen3-max-thinking kimi-k2.5 Avg.
RM-NLHF Win 57.70 54.96 49.47 47.19 52.33
Outcome-Only Win 34.40 37.67 43.98 41.58 39.41
Tie 7.90 7.37 6.55 11.23 8.26
Figure 7: Mean response length during downstream RL training.
Table 5: Evaluation results on downstream tasks through test-time scaling.
@@ 4.4 Performance of Downstream Tasks through Reinforcement Learning
4.5 Performance of Downstream Tasks through Test-time Scaling
@@ G.1 Benchmarks
RM-Bench (Liu et al., 2024c) 2024.10 Chat, Safety, Code, Math Verifiable correctness, deliberate errors
@@
SCAN-HPD (Wang et al., 2025d) 2025.05 Writing, Roleplay, Knowledge, Coding, Math, Reasoning Majority human voting
HelpSteer3 (Wang et al., 2025c) 2025.05 General chat, STEM, Code, Multilingual Majority human voting
@@ G.2 Training Datasets
HelpSteer3 (Wang et al., 2025c) 36,255 General Domain
@@ H.1 Ablation Study on MetaRM
Appendix I Understanding Reward Modeling from Theoretical Perspective: From Statistical Correlation to Causal Alignment
I.1 Notation and Idealized Reward Modeling
We consider the classical pairwise preference reward task: given an input prompt $x$, the model generates two candidate responses $y_{A},y_{B}\in\mathcal{Y}$, and human annotators provide a preference label $l\in\{A,B\}$.
For any triplet $(x,y_{A},y_{B})$, we assume there exists an ideal evaluation dimension vector $\mathbf{z}=[z_{1},z_{2},\dots,z_{K}]^{\top}$ and weight vector $\mathbf{w}=[w_{1},w_{2},\dots,w_{K}]^{\top}$, where each dimension $z_{k}$ characterizes an interpretable quality attribute (e.g., factual accuracy, logical coherence, linguistic politeness, response length, etc). Weight $w_{k}\geq 0$ captures the importance of that dimension in the current comparison scenario. The combination of dimensions and weights constitutes what we refer to as the reward causality. For generative reward models, reward causality is sometimes contained in the critique rather than direct numerical computation.
Note: Both the dimension and weight vector $\mathbf{w}$ are context-dependent—only after given a specific $(x,y_{A},y_{B})$ can we infer which dimensions actually participate in human judgment and their relative importance. Ideally, the weights of key dimensions should approach 1, while irrelevant dimensions should have weights approaching 0. For example: in knowledge-intensive QA tasks, the weight of "factual accuracy" far exceeds "linguistic politeness"—a factually incorrect response, no matter how politely expressed, typically cannot prevail over a factually accurate but plain-toned response.
I.2 Summary of Reward Modeling Mechanisms
Since the emergence of traditional BT reward modeling, numerous improvement schemes have emerged to address the limitations of reward modeling. Their common approach is to enhance model performance through causal alignment.
This article formalizes "causality" as a combined representation of evaluation dimensions and weights, and systematically reviews the following schemes shown in Table 12.
Table 12: Summary of Different Reward Modeling Mechanisms
Method Modeling Approach Modeling Formula Learning Objective
Bradley-Terry Reward Modeling Compress high-dimensional evaluation space into a single scalar score $P_{\phi}(y_{A}\succ y_{B}|x)=\sigma\left(r_{\phi}(x,y_{A})-r_{\phi}(x,y_{B})\right)$ Learn scalar utility function through minimizing negative log-likelihood loss, such that $r_{\phi}(x,y)\approx\mathbf{w}^{\top}\mathbf{z}$
Conditional BT Reward Modeling Explicitly model causality as part of the input $P_{\phi}(y_{A}\succ y_{B}|x,\mathbf{d},\mathbf{w})=\sigma\left(\mathbf{w}^{\top}r_{\phi}(x,y_{A}|\mathbf{d})-\mathbf{w}^{\top}r_{\phi}(x,y_{B}|\mathbf{d})\right)$ Learn decomposed rewards conditioned on dimensions, based on predefined dimension description vector $\mathbf{d}$ and weight vector $\mathbf{w}$
Generative Reward Modeling with Outcome Supervision Explicitly model causality as part of the output $P_{\theta}(c,l|x,y_{A},y_{B})=P_{\theta}(c|x,y_{A},y_{B})\cdot\underbrace{P_{\theta}(l|x,y_{A},y_{B},c)}_{\text{supervised}}$ Generate critique $c$ and preference label $l$, using only preference label correctness as supervision signal, expecting causal capability through result correctness constraint
Generative Reward Modeling with Process Supervision Explicitly supervise the causality in the output $P_{\theta}(c,l|x,y_{A},y_{B})=\underbrace{P_{\theta}(c|x,y_{A},y_{B})}_{\text{supervised}}\cdot\underbrace{P_{\theta}(l|x,y_{A},y_{B},c)}_{\text{supervised}}$ Directly optimize causal alignment for explicit critique supervision
I.3 Bradley-Terry Reward Modeling
Modeling Mechanism: Traditional reward models assume a scalar utility function $r_{\phi}(x,y)$. The core idea is to compress the high-dimensional evaluation space $\mathbf{z}$ into a single scalar score. The preference probability follows the Bradley-Terry (BT) model:
$P_{\phi}(y_{A}\succ y_{B}|x)=\sigma\left(r_{\phi}(x,y_{A})-r_{\phi}(x,y_{B})\right)$ (9)
The model is trained by minimizing the negative log-likelihood loss $\mathcal{L}_{\text{BT}}$. Ideally, the scalar reward should approximate the inner product of dimension features and true weights:
$r_{\phi}(x,y)\approx\mathbf{w}^{\top}\mathbf{z}$ (10)
Failure Modes: The BT model is fundamentally a statistical learning approach that estimates preference patterns from empirical data distributions. While effective at the dataset level, this statistical nature introduces two primary sources of bias that compromise its discriminative power on individual samples:
1. Sample-Level Causal Ambiguity: The BT model suffers from causal sparsity of supervision: for a sample $(x,y_{A},y_{B},l)$, the training objective only models the preference label $l$ without capturing which dimensions determined the preference and their relative importance (Wang et al., 2024a; Minghao Yang, 2024). To compensate, the model relies on statistical aggregation extracted from patterns of similar samples in the training data. Learning from aggregated statistics rather than sample-specific criteria leads to inaccurate causal estimation at the individual sample level.
2. Dataset-Level Causal Ambiguity: Beyond similar samples, global dataset patterns also introduce bias (Liu et al., 2024b). Since the BT model uses global empirical risk minimization for optimization, if a large proportion of "better" responses in the dataset share features like "high detail" or "polite tone," the model may increase their weights globally, even when irrelevant to specific samples.
3. Lack of Reasoning Capability: Without reasoning capability, the BT model tends to capture surface-level correlations with label $l$, failing to model dimensions requiring deep reasoning (Chen et al., 2025; Guo et al., 2025c; Liu et al., 2025). For example, in math tasks, it struggles to learn "answer correctness" (which needs symbolic verification) and instead relies on proxy features like "solution detail."
In summary, the traditional BT model introduces causal ambiguity (or spurious correlations) for individual samples and lacks reasoning capability. It is only suitable for dimensions that are globally important and surface-observable (e.g., style, politeness, detail).
I.4 Conditional BT Reward Modeling
Modeling Mechanism: To alleviate the causal ambiguity of BT reward modeling, recent work Wang et al. (2025b) proposes explicitly modeling evaluation dimensions into the reward model. Although existing schemes mostly focus on single-dimension scenarios (implicitly assuming the weight of this dimension is 1), they can naturally extend to more general settings:
$P_{\phi}(y_{A}\succ y_{B}|x,\mathbf{d},\mathbf{w})=\sigma\left(\mathbf{w}^{\top}r_{\phi}(x,y_{A}|\mathbf{d})-\mathbf{w}^{\top}r_{\phi}(x,y_{B}|\mathbf{d})\right)$ (11)
where $\mathbf{d}=[d_{1},d_{2},\dots,d_{K}]$ is a predefined dimension description vector, $r_{\phi}(x,y|\mathbf{d})\in\mathbb{R}^{K}$ is the decomposed reward conditioned on dimensions, and $\mathbf{w}$ is a manually specified weight vector.
Failure Modes: This mechanism directly addresses causal ambiguity by pre-specifying causal auxiliary evaluation. Several works Gu et al. (2026); Wang et al. (2025b); Yu et al. (2025c); Wang et al. (2025d) leverage LLM’s reasoning capability for conditional generative reward modeling. Note that while these works address the lack of reasoning capability in BT models, all conditional reward modeling mechanisms face a significant limitation.
Specifically, since the dimension set $\mathbf{d}$ must be predefined, this method is only effective when key dimensions are highly determined. When facing (1) open-domain tasks with high openness such as open-domain dialogue and role-playing, and (2) tasks requiring complex and diverse process rewards such as multi-turn task planning and long-horizon reasoning, evaluation dimensions and weights are typically difficult to pre-specify, or even if pre-specified, are difficult to maintain consistency across all samples. Conditional reward modeling fails in these scenarios.
In summary, pre-specifying evaluation dimensions and weights endows BT models with the ability to incorporate causal reasoning for reward assignment. Compared to traditional BT models, they are more resistant to spurious correlations, but their applicability is narrow due to the need to pre-specify dimensions and weights. They are more suitable for tasks with clear rules, such as "syntax correctness" and "query efficiency" in SQL generation, or "citation reliability" and "argumentation completeness" in fact-checking.
I.5 Generative Reward Modeling with Outcome Supervision
Modeling Mechanism: Generative reward modeling changes the paradigm of reward modeling. Unlike conditional reward modeling which models causality in the input part, it models causality in the output part. Given input $(x,y_{A},y_{B})$, the model is required to generate a text sequence containing "critique" and "outcome". Some methods require the model to explicitly output evaluation dimensions and weights in the critique before giving the outcome (Liu et al., 2025), while others implicitly include this process in the critique without explicitly outputting related content (Wang et al., 2025c; Chen et al., 2025; Guo et al., 2025c; Huang et al., 2025; Xu et al., 2025a).
$P_{\theta}(c,l|x,y_{A},y_{B})=P_{\theta}(c|x,y_{A},y_{B})\cdot\underbrace{P_{\theta}(l|x,y_{A},y_{B},c)}_{\text{supervised}}$ (12)
where $c$ is a natural language analysis of the superiority of the two responses (i.e., critique), which explicitly points out the evaluation dimensions $\mathbf{z}$ and local weights $\mathbf{w}$ in the semantic space. $l\in\{A,B\}$ is the final preference label derived from the critique. Training typically uses correctness of the preference label as a supervision signal to optimize the probability of the model predicting the correct label $l$.
Failure Modes: Compared to conditional reward modeling, GRM has the following advantages: (1) Automatic Causal Discovery: By generating critique $c$, the model is forced to explicitly identify key dimensions of the current sample (e.g., code logic is more important than comments for this problem), achieving identification of key dimensions and dynamic weight allocation. Additionally, it can use test-time scaling to find better dimensions and weights. (2) Resistance to Global Ambiguity: Since GRM requires explicit causal analysis and learning for each sample, it is harder to be affected by global statistical biases.
Despite the great potential of GRM, our work identifies a critical misconception in existing approaches: they actually expect the model to learn good causal capability through outcome correctness. However, we find this introduces significant noise to training as discussed above.
I.6 Generative Reward Modeling with Process Supervision
To directly align reward modeling with causality, we introduce process supervision, i.e., supervision of critique. Our paradigm requires GRM to optimize the correctness of critiques. The modeling approach remains unchanged, but the supervision signal changes:
$P_{\theta}(c,l|x,y_{A},y_{B})=\underbrace{P_{\theta}(c|x,y_{A},y_{B})}_{\text{supervised}}\cdot\underbrace{P_{\theta}(l|x,y_{A},y_{B},c)}_{\text{supervised}}$ (13)
Important evidence supporting this argument comes from the difference in critique word frequency distributions of models trained with the two supervision methods (see Section 4.9 and Table 7 for details).
I.7 Theory-Practice Gap
Although our theoretical framework indicates that we should conduct reward modeling by supervising critique, there are several challenges in practice:
Cost of High-Quality Human Critiques: Currently, our approach to critique supervision relies on high-quality human critiques; however, collecting such data is both difficult and resource-intensive.
Difficulty in Obtaining Accurate Supervision: We use the similarity between GRM-generated and human critiques as the supervision metric, achieving 85.71% accuracy. This imperfect signal leads to: (1) suboptimal reward model performance, and (2) training instability. Fortunately, we have addressed this issue by introducing outcome regularization, which enables stable long-term training and ultimately achieves optimal performance.
Despite these theory-practice gaps, we believe our work points to a promising direction for future reward modeling, i.e., through substantial engineering efforts to obtain higher-quality human critiques, thereby obtaining high-quality critique supervision signals and ultimately more accurate reward models. Additionally, using human critique supervision may not be necessary, but this requires further research.
Appendix J Prompt